OpenStudy (anonymous):

HELP PLEASE?!?!?!?! :) Find the solution set of |2x-3| + 1 = 6. A. x = 1 or x = -1 B. x = 5 or x = -2 C. x = 4 or x = -1 D. x = 4 or x = -4

jimthompson5910 (jim_thompson5910):

first you need to isolate |2x-3| to do so, you subtract 1 from both sides

OpenStudy (ageta):

C. x = 4 or x = -1

jimthompson5910 (jim_thompson5910):

|2x-3| + 1 = 6 |2x-3| + 1-1 = 6-1 |2x-3| = 5 making sense so far?

OpenStudy (ageta):

Let's solve your equation step-by-step. |2x−3|+1=6 Step 1: Add -1 to both sides. |2x−3|+1+−1=6+−1 |2x−3|=5 Step 2: Solve Absolute Value. |2x−3|=5

OpenStudy (ageta):

We know either 2x−3=5 or 2x−3=−5

OpenStudy (ageta):

x−3=5(Possibility 1) 2x−3+3=5+3(Add 3 to both sides) 2x=8 2x/2=8/2 (Divide both sides by 2) x=4/2x−3=−5(Possibility 2) 2x−3+3=−5+3(Add 3 to both sides) 2x=−2/2x 2=−2/2 (Divide both sides by 2) x=−1 Answer: x=4 or x=−1